OpenStudy (anonymous):

What is the d value in the following arithmetic sequence? -7, -2, 3, 8, ...

OpenStudy (anonymous):

What are the next four terms in the following arithmetic sequence? -7, -2, 3, 8, ...

OpenStudy (anonymous):

What is the 25 the term in the following arithmetic sequence? -7, -2, 3, 8, ...

OpenStudy (anonymous):

Find the four arithmetic means between -21 and -36.

OpenStudy (texaschic101):

-7,-2,3,8,13,18,23,28 d = 5 what is the 25....do you mean the 25th term ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

an = a1 + ((n - 1)d d = 5 n = 25 a1 = -7 a25 = -7 + (25 - 1)5 a25 = - 7 + 120 a25 = 113 the 25th term is 113

OpenStudy (anonymous):

Find the four arithmetic means between -21 and -36.

OpenStudy (anonymous):

1. Find S25 for -7+ -2 + 3 + 8 +…Type the numeric answer only. 2. Find S25 for 1/2+1+3/2+2+...Enter your answer as a fraction. Do not type any spaces. Answer 3. Find S10 for 2+5+8...Type the numeric answer only. 4. Find S10 for -1+-7+-13+-19...Type the numeric answer only. Answer Thanks!!!

OpenStudy (texaschic101):

4 arithmetic means between -21 and - 36 -21,____, ____ ,_____, _____, -36 -21 + 5d = - 36 5d = -36 + 21 5d = - 15 d = -3 -21 +(-3) = -21 - 3 = - 24 <--2nd number -21 + 2(-3) = -21 - 6 = - 27 <-- 3rd number -21 + 3(-3) = -21 - 9 = - 30 < -- 4th number -21 + 4(-3) = - 21 - 12 = - 33 < -- 5th number

OpenStudy (texaschic101):

I have already found the 25th term for the sequence -7,-2,3,8, and it was 113 find 25th term for the sequence 1/2,1,3/2,2 d = 1/2 n = 25 a1 = 1/2 an = a1 + (n - 1)d a25 = 1/2 + (25 - 1)1/2 a25 = 1/2 + 24/2 = 25/2 The 25th term is 25/2 find 10th term for sequence 2,5,8 d = 3 a1 = 2 n = 10 an = a1 + (n - 1)d a10 = 2 + (25 - 1)3 a10 = 2 + 24/3 a10 = 2 + 6 = 8 The 10th term is 8 Find 10th term for -1,-7,-13,-19 d = -6 a1 = -1 n = 10 an = a1 + (n - 1)-6 a10 = -1 + (10 - 1)-6 a10 = -1 - 54 a10 = - 55 The 10th term is - 55

OpenStudy (texaschic101):

wait....I must of messed up on the 3rd one... find the 10th term of sequence 2,5,8 d = 3 a1 = 2 n = 10 an = a1 + (n - 1)3 a10 = 2 + (10 - 1)3 a10 = 2 + 27 a10 = 29 The tenth term is 29 sorry about that